MILLIMETER-WAVE SPECTROSCOPY OF AMINOMALONONITRILE

\underline{ROMAN A. MOTIYENKO}, LAURENT MARGULÈS, Laboratoire PhLAM, UMR 8523 CNRS - Université Lille 1, 59655 Villeneuve d’Ascq Cedex, France; JEAN-CLAUDE GUILLEMIN, Institut des Sciences Chimiques de Rennes, UMR 6226 CNRS - ENSCR, 35708 Rennes Cedex 7, France.

In low-temperature conditions of the interstellar medium pure HCN and mixtures of HCN with H_{2}O and NH_{3} subjected to high-energy sources can be converted into different HCN polymers^{a}. In this context the HCN trimer – aminomalononitrile (AMN, H_{2}NCH(CN)_{2}) is an interesting candidate for astrophysical detection. We measured the rotational spectrum of AMN using the Lille BWO-based fast scan spectrometer (120 – 180 GHz) and the spectrometer based on solid state sources (225 – 250 GHz). The spectroscopic work was supported by high level ab initio calculations. The spectra observed were assigned only to the asymmetric conformer. The first analysis revealed the regular doublet splittings of rotational lines indicating the possibility of large-amplitude motion (LAM) of amino group. Further examination of the spectra revealed the existence of additional splittings most probably due to the second LAM of amino group. Since the equilibrium structure of AMN has no symmetry, in the group theoretical formalism these two LAMs can be described using G_{4} permutation-inversion group. Despite the complexity of the LAMs all the assigned rotational transitions were fitted within experimental accuracy using Pickett’s SPFIT program. As a result, a reliable dataset for astrophysical observations was provided. This work is supported by the french program ”Physique et Chimie du Milieu Interstellaire (PCMI, INSU-CNRS)” and the Centre National d’Etudes Spatiales (CNES).

aGerakines, P.A., Moore, M.H, Hudson, R.L., Icarus, 2004, 170, 203